Einstein throws a ball from the ground to the top of the goal. If it returns to the ground after 7 seconds, what is the height of the goal?
1
Expert's answer
2020-10-12T07:44:06-0400

The ball starts falling from the top of the goal with zero initial velocity. Thus, the distance travelled by ball (height of the goal) is:


"h = \dfrac{gt^2}{2}"

where "g = 9.81m\/s^2" is the gravitational acceleration and "t = 7s" is the time of falling. Thus, obtain:


"h = \dfrac{gt^2}{2} = \dfrac{9.81\cdot 7^2}{2} \approx 240.345m"

Answer. 240.345 m.